OpenStudy (anonymous):

What is the value of the expression 8–^2?

rvc (rvc):

is it \(\huge 8^{-2}\)

OpenStudy (anonymous):

yes

rvc (rvc):

here we apply the law of indices \(\Large\color{red}{a^{-n}=\frac{1}{a^n}}\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

It's saying \[8^{-2}=\frac{ 1 }{ 8^{2} }\]

OpenStudy (anonymous):

is it 1 _ 16

OpenStudy (anonymous):

\[\frac{ 1 }{16? }\]

OpenStudy (anonymous):

\[8^{2}=8*8\]

OpenStudy (anonymous):

\[\frac{ 1 }{64 }\]

rvc (rvc):

correct :)

OpenStudy (anonymous):

Yes, that would be it.